### v2

# Manolis Antonoyiannakis American Physical Society, 1 Physics Ellipse, College Park, MD 20740

Abstract
The sensitivity of Impact Factors (IFs) to journal size causes systematic bias in IF rankings, in a process akin to stacking the cards: A random “journal” of $n$ papers can attain a range of IF values that decreases rapidly with size, as $\sim 1/\sqrt{n}$ The Central Limit Theorem, which underlies this effect, allows us also to correct for it by standardizing citation averages for scale and subject in a geometrically intuitive manner analogous to calculating the $z$-score. We thus propose the $\Phi$ index, a standardized scale- and subject-independent citation average. The $\Phi$ index passes the “random sample test”, a simple check for scale and subject independence that we argue ought to be used for every citation indicator. We present $\Phi$ index rankings for 12,173 journals using data from the 2020 Journal Citation Reports. We show how scale standardization alone affects rankings, then we demonstrate the additional effect of subject standardization for monodisciplinary journals, and we discuss how to treat multidisciplinary journals. $\Phi$ index rankings offer a clear improvement over IF rankings. And because the $\Phi$ index methodology is general, it can also be applied to compare individual researchers, universities, or countries.
keywords:
Science of Science , Impact Factor , Citation Distributions , Bibliostatistics , $\Phi$ index
journal: arXiv
url]www.bibliostatistics.org
1 Why a scale-independent citation indicator?
1.1 Impact Factors are scale-dependent
Every year at the end of June, when the Journal Impact Factors are announced by Clarivate Analytics, a familiar pattern occurs (Antonoyiannakis, 2018): (a) The high-Impact-Factor ranks are always populated by small journals that typically publish no more than a few hundred papers per discipline per year. For example, all the top 50 IF ranks—and all but 13 of the top 51–250 ranks—of the 2020 Journal Citation Reports (JCR) are filled by journals that publish fewer than 1000 papers/year. Of the 13 slots in the ranks 51–250 that are not occupied by small journals, we find 12 mid-sized journals (between 1000-5000 papers/year), and only one large journal ($>5000$ papers/year), in the 196th rank. (b) Small journals also occupy the low-Impact-Factor ranks—for example, all but one of the bottom 1000 ranks in the 2020 JCR, the one exception being a mid-sized journal. (c) Large journals that publish more than 5000 papers/year have modest Impact Factors that are neither very small nor very large, and tend to occupy unremarkable positions in the ranks—for example, there are 10 large journals with IF ranging from 2.5–14 and rank positions from 196 to 5189. Thus, large journals are precluded from reaching high IF ranks but are also “protected” from falling to the lowest of ranks as well. (d) Single papers affect the Impact Factors of hundreds—sometimes thousands—of small journals. For example, in 2017, the IF of 818 journals increased by more than 25% by their single top-cited paper, while one in 10 journals had their IF boosted by more than 50% by their top three cited papers (Antonoyiannakis, 2020).
Clearly, therefore, journal size affects IF rankings. This happens because despite the averaging process entailed in the IF calculation, IFs are not size-independent. They are size-normalized but size-dependent, because they are drawn from populations with high variance in citations, as explained by the Central Limit Theorem (Antonoyiannakis, 2018). Remarkably, and despite this recurrent size-dependent behavior of Impact Factors every year, “the overwhelming majority of researchers, bibliometricians, administrators, publishers, and editors continue to use them without realizing or acknowledging the problem.” (Antonoyiannakis, 2020).
To the extent that IF rankings are important to the research community, high-IF journals have a perverse incentive to stay small in order to remain competitive in IF rankings (Antonoyiannakis, 2019; Antonoyiannakis & Mitra, 2009). The editorial choice of the vast majority of high-IF journals to publish in the low-hundreds of research papers per year in any given research area may thus be influenced more by the mechanics of IF sensitivity to journal scale than by a genuine desire to stay small, or an organic response to rapidly developing fields and the needs of a dynamic and diverse audience of researchers (Campbell, 2008).
What can be done to remedy this situation and level the playing field? One approach would be to abandon citation averages (and thus Impact Factors) altogether in assessing research journals. Indeed, several other flaws of IF rankings have long been identified, such as the skewness of citation distributions, the subject dependence of citation averages, the numerator/denominator asymmetry, journal self-citations, the length of citation window, IF inflation, IF gaming and engineering, or the document-type citation disparity (Amin & Mabe, 2004; Moed, 2005; Larivière & Sugimoto, 2019; Pulverer, 2013; Garfield, 1996; Gingras, 2016; Wouters et al., 2019; Siler & Larivière, 2022). Growing concern for these shortcomings and their repercussions in IF rankings have led to public calls for abandoning, de-emphasizing, or at least complementing IFs with more reliable metrics (San Francisco Declaration on Research Assessment, 2012; Gaind, 2018) and detailed, statistically sensible visualizations (Adams et al., 2019).
In our judgment, Impact Factors are currently too entrenched in the minds of scientists and administrators to completely do away with them. Besides, the concept of an average is appealing for its simplicity. A more sensible approach may be therefore to provide a technical solution that corrects the scale-dependence of citation averages—in the same fashion that the subject-dependence of IFs has been addressed by indicators like SNIP (Waltman, 2013) or the Journal Citation Indicator (Clarivate, 2021; Szomszor, 2021)—and argue for its acceptance by the community. This is the main aim of our paper.
Fortunately, such a solution is simple and can be found with the help of the Central Limit Theorem. Our methodology allows us to also correct for the subject dependence of citation averages, in an elegant and statistically sensible manner that is also distinct from other approaches.
In this paper, we introduce the concept of standardized citation averages. First, we perform a numerical experiment, where we show how a random journal—formed by consecutive unbiased samplings of actual papers in a certain subject—has a citation average whose range of values clearly depends on its size, hence the need for a correction. We call this experiment the “random sample test.” Then, we express the Central Limit Theorem algebraically, thus bringing out the standardization of the citation average to correct for the scale dependence, which leads us to introduce the $\Phi$ index. We present an elegant geometric interpretation of the $\Phi$ index, followed by a more formal interpretation as a $z$-score. We show how it accounts for both scale- and subject-dependence of citation averages, thus passing the random sample test. We then present $\Phi$ index data for journals, using data from the 2020 JCR.
Our work is aligned with the recommendations of scrutinizing and updating metrics (Hicks et al., 2015), adding to the variety of journal-based metrics and accounting for article-type citation disparity and subject-dependent citation practices (San Francisco Declaration on Research Assessment, 2012; Moed et al., 2012).
1.2 A randomly formed journal reveals the problem
Consider a population of all papers published in a two-year window for which we collect their citations during the following (third) year. (We use these time windows for publications and citations merely for relevance to Impact Factors, but our analysis applies to averaging over any time windows.) Let us call the population’s global citation average $\mu$ and its citation standard deviation $\sigma$. For example, the 3,537,118 papers (articles and reviews) published in 2018–2019 have $\mu$=4.11 and $\sigma$=12.5 (2020 JCR).
Now imagine we draw thousands of random samples of size $n$ from this population and calculate their citation average, $f_{n}$, which is essentially the Impact Factor of an $n$-sized “random journal.” We form a scatter plot of the values ($n,f_{n}$) as shown in Fig. 1. As we can see, smaller journals have a better chance of scoring a high IF (but also low IF); while as journal size increases, variations from the global mean get smaller. Clearly, the playing field is not level.
Let us understand this effect quantitatively. Under random sampling conditions, the sampling distribution of the sample means is normally distributed, so that
$\mu-k\sigma_{n}\leq f_{n}\leq\mu+k\sigma_{n}\;,\qquad(k=1,2,3,...),$ (1)
where $k$ is an integer that increases with statistical accuracy, e.g., when $k=1$ then Eq. (1) is expected to capture 68% of observations, $k=2$ for 95%, $k=3$ for 99% of observations, etc. The Central Limit Theorem (De Veaux, Velleman, & Bock, 2014) allows us to express the sample standard deviation $\sigma_{n}$ in terms of the population standard deviation $\sigma$ as follows,
$\sigma_{n}=\frac{\sigma}{\sqrt{n}}.$ (2)
Equations (1) and (2) imply that
$\mu-\frac{k\sigma}{\sqrt{n}}\leq f_{n}\leq\mu+\frac{k\sigma}{\sqrt{n}}\;,\qquad\text{where}\;\;\begin{cases}k=1,&\text{for 68\% of random journals}\\ k=2,&\text{for 95\% of random journals}\\ k=3,&\text{for 99\% of random journals}\end{cases}$ (3)
or
$-\frac{k\sigma}{\sqrt{n}}\leq f_{n}-\mu\leq+\frac{k\sigma}{\sqrt{n}}\;,$ (4)
which can be rewritten as
[…]
where $\text{max}(\Delta f_{n})\equiv\text{max}(|f_{n}-\mu|)$ is the maximum distance (or range) of $f_{n}$ values from the global mean $\mu$. Equation (5) says that for randomly formed journals, the range of $f_{n}$ values from the global mean $\mu$ is inversely proportional to the square root of the journal size. Herein lies the scale dependence of citation averages.
Figure 1: Dependence of citation average $f_{n}$ on journal size $n$ for a random journal. Each dot represents $f_{n}$ for a random selection of $n$ papers within the field of cell biology. The range of $f_{n}$ values is centered on the global citation average $\mu$, while the upper and lower envelopes are proportional to $k\sigma/\sqrt{n}$ as shown in Eqs. (3, 4)
While Eq. (5) applies to randomly formed journals, it turns out that 78% of all actual journals (9571 out of 12173 journals in the 2020 JCR) have IFs within the range predicted by the Central Limit Theorem (with $k=3$). So the Central Limit Theorem is directly relevant to journals in practice. It is not just of academic interest but allows us to predict and understand citation averages of real journals. See Figures 2, 3.
Figure 2: Left panel: Dependence of citation average $f_{n}$ on journal size $n$ for all 11639 journals in the 2017 JCR. Shown in blue and orange are the upper and lower limits, respectively, predicted by the Central Limit Theorem for random journals. 73% of all journals lie within these bounds. Right panel: Similar plot for the field of biochemistry and molecular biology. 65% of all journals lie within the CLT bounds.
Figure 3: Left panel: Dependence of citation average $f_{n}$ on journal size $n$ for all business journals in the 2017 JCR. Shown in blue and orange are the upper and lower limits, respectively, predicted by the Central Limit Theorem for random journals. 66% of all journals lie within these bounds. Right panel: Similar plot for the field of biochemistry and molecular biology. 63% of all journals lie within the CLT bounds.
1.3 Standardization of citation averages corrects their scale dependence
The Central Limit Theorem allows us to standardize citation averages and remove their scale dependence. Here is how. Upon rewriting Eq. (3) as
$-k\leq\frac{(f_{n}-\mu)\sqrt{n}}{\sigma}\leq k\;,\qquad\begin{cases}k=1&\text{for 68\% of random journals}\\ k=2&\text{for 95\% of random journals}\\ k=3&\text{for 99\% of random journals}\end{cases}$ (6)
we realize that the quantity
$\Phi\equiv\frac{(f_{n}-\mu)\sqrt{n}}{\sigma},$ (7)
which we hereby call the $\Phi$ index, is equal to the distance of the IF from the mean $(f_{n}-\mu)$, normalized to the sample (i.e., journal) standard deviation $(\sigma/\sqrt{n})$ from Eq. (2). The expression in Eq. (7) may be familiar to some readers from the statistics of standardizing averages (De Veaux, Velleman, & Bock, 2014, pp. 110–118). Indeed, the $\Phi$ index is a standardized average analogous to the $z$-score in statistics, which measures how many standard deviations a data point is away from the mean. Equation (6) shows that the $\Phi$ index is bounded by a constant (integer $k$), unlike the citation average $f_{n}$ that has size-dependent bounds, as shown in Eq. (3). This is a first indication that the $\Phi$ index is size-independent; we will shortly provide a numerical demonstration as well.
We can calculate a $\Phi$ index for any sample drawn from a distribution of research papers with a well-defined ($\mu,\sigma$). Therefore, we can also use it for standardizing citation averages across different research fields. To compare $\Phi$ indexes of journals in different fields, all we need to know are the citation mean ($\mu$) and standard deviation ($\sigma$) of the papers in each field.
Beyond the analogy with the $z$-score in statistics, the $\Phi$ index has also an elegant geometric interpretation, as we explain below.
1.4 A geometric interpretation of the $\Phi$ index
Let us plot of the citation average $f_{n}$ as a function of journal size $n$ for various randomly formed journals that are drawn from a distribution of research papers described by ($\mu,\sigma$). We wish to compare two different-size journals represented by the points $A$ and $B$ in the plot, with respective coordinates ($f_{1},n_{1}$) and ($f_{2},n_{2}$).
Consider the journal with the maximum citation average at size $n_{1}$, namely, $f_{1,max}=\mu+k\sigma/\sqrt{n_{1}}$ from Eq. (3), which we denote with the point $A^{\prime}$. The journals $A$ and $A^{\prime}$ differ in selectivity: Journal $A^{\prime}$ is the most selective journal possible at size $n_{1}$, because it contains all the highest-cited papers one can find for a journal of this size. A journal with a different size $n_{2}$ will have a different maximum citation average as dictated by Eq. (3), namely, $f_{2,max}=\mu+k\sigma/\sqrt{n_{2}}$; we denote this journal with point $B^{\prime}$ in the plot. The points $A^{\prime}$ and $B^{\prime}$ are equivalent: They represent the highest possible citation average for journals of different sizes, and correspond to a paper selection process that is maximally selective at each size.
We posit that journals $A$ and $B$ will be equivalent if they have the same score in their respective scales, which are size-dependent because the maximum scores attainable to different-sized journals are different. As we can see from the geometrical construction in Fig. 4, points $A$ and $B$ are equivalent citation-wise if they are equidistant from the global citation mean of their field $\mu$, where distances are normalized to the IF range at any given size. In geometrical terms, $A$ and $B$ are equivalent if
$\frac{\Delta f_{1}}{\text{max}(\Delta f_{1})}=\frac{\Delta f_{2}}{\text{max}(\Delta f_{2})}$ (8)
and therefore
$\frac{(f_{1}-\mu)\sqrt{n_{1}}}{\sigma}=\frac{(f_{2}-\mu)\sqrt{n_{2}}}{\sigma}$ (9)
or
$\Phi_{1}=\Phi_{2}.$ (10)
In other words, journals $A$ and $B$ will be equivalent if they are equally selective for their size. This geometric interpretation provides an intuitive understanding of the $\Phi$ index.
Figure 4: Geometric interpretation of the $\Phi$ index.
1.5 The $\Phi$ index as a $z$ score.
The geometric interpretation of the $\Phi$ index agrees intuitively with our expectation that journals A and B in Fig. 4 ought to be equivalent in terms of their citation accrual. But it also has a more rigorous interpretation, as the $z$ score from statistics. It tells us how far from the average a value is, where the distance from the average is normalized to the standard deviation of the population of $n-$sized journals. Thus, a $z$ score of 1 tells us that our score is 1 standard deviation away from the mean. Interpreted as a $z$ score, the $\Phi$ index is a standardized metric and thus allows also comparison of scores (impact factors) that are measured on different scales (different $\mu,\sigma$) in analogy with the use of standardized averages to compare athletes’ performances across different sports or students’ scores across different academic disciplines (De Veaux, Velleman, & Bock, 2014, pp. 111–112).
In the most general form, we have proposed (Antonoyiannakis, 2018) the $\Phi$ index as
$\Phi=\frac{f-\mu_{s}}{\sigma_{s}/\sqrt{N_{2Y}}},$ (11)
where $\mu_{s},\sigma_{s}$ are the global average and standard deviation of the citation distribution of all papers in the subject of the journal in question. Here, we denoted with $f$ and $N_{2Y}$ the journal’s citation average (IF) and biennial size, respectively. Also, we have used $k=1$ for simplicity and without loss of generality.
2 A standardized citation indicator accounts for scale and subject
To calculate a journal’s $\Phi$ index from Eq. (11), we need to know the quantities $\mu_{s}$ and $\sigma_{s}$ for the journal’s research subject. For example, if we were to treat all 3,537,118 papers published in all journals in 2018–2019 as if they form a single subject, we would have $\mu=4.11,\sigma=12.5$. A comparison of the $\Phi$ index of all journals in this case would be free from the size-dependence bias of citation averages, and in this regard would be fairer than a comparison of Impact Factors.
But we can do better. Citation indicator comparisons need to account for differences in citation practices across research fields. A glance at Eq. (11) shows how we can readily use the $\Phi$ index to standardize both for size and subject. To compare journals that publish in different research subjects, all we need is find the $\mu_{s}$ and $\sigma_{s}$ of each subject, and use them to calculate the $\Phi$ index of each journal. A comparison of subject-specific $\Phi$ indexes would account for size and subject differences. In Table 1 we show different ($\mu_{s},\sigma_{s}$) for several subject categories in the 2020 Journal Citation Reports (JCR) of Clarivate Analytics. As we can see, they each span more than one order of magnitude, $\mu_{s}$ ranging from 0.5–8 and $\sigma_{s}$ ranging from 1–50.
2.1 What is a subject?
Equation (11) says that a journal’s standardized citation average depends on its community of journals, all of which define the subject and determine the global quantities $\mu_{s}$ and $\sigma_{s}$, which are computed for all papers in the subject. The notion that a journal’s score is not defined in absolute terms but in relation to its environment is key to Eq. (11). Indeed, what difference does it make if a journal’s citation score is 1, 10, or 100? In itself, the number can mean anything; it is in connection to its community of peer journals that the number takes meaning. This is the logic behind the standardized average, and we now see why the $\Phi$ index accounts naturally for subject standardization, in addition to standardization of size.
But if a journal is measured in relation to its peers, then the choice of peers affects the journal’s score by definition. So the selection of the journals that form a subject is important. Clearly, if we include in a subject those journals with higher or lower citation averages, we push the subject’s global average $\mu_{s}$ to correspondingly higher or lower values. The value of $\mu_{s}$ affects in turn the sign of $(f-\mu_{s})$, which is multiplied by a size-sensitive factor $\sqrt{N_{2Y}}$ in the numerator of the $\Phi$ index. Thus whenever the addition or removal of journals in a subject causes the term $(f-\mu_{s})$ to flip sign, the corresponding journal’s $\Phi$ index can jump abruptly in the rankings, particularly for large journals. This caveat is important to keep in mind, especially if we try to over-analyze subjects by describing them into smaller journal lists. At any rate, the decision of which journals to include in each subject category has important ramifications in subject-based citation rankings. But this issue is hardly unique to the $\Phi$ index rankings.
In this paper, we rely on the 229 subject categories in the Clarivate Analytics 2020 JCR.
Field (subject) $\mu_{s}$ $\sigma_{s}$ no. papers no. journals
ALL 4.11 12.50 3,537,118 12173
AREA STUDIES 1.09 2.01 5301 80
ASTRONOMY & ASTROPHYSICS 4.53 14.80 42725 60
BIOCHEMISTRY & MOLECULAR BIOLOGY 5.65 18.12 118555 279
CARDIAC & CARDIOVASCULAR SYSTEMS 4.77 15.85 40771 142
CELL BIOLOGY 7.20 14.84 61131 191
CHEMISTRY, MULTIDISCIPLINARY 6.86 12.93 156779 171
COMPUTER SCIENCE, CYBERNETICS 5.53 17.14 3210 22
CULTURAL STUDIES 1.00 2.231 3167 45
ENERGY & FUELS 7.23 11.72 85899 108
ENGINEERING, ENVIRONMENTAL 8.07 10.64 40179 54
EVOLUTIONARY BIOLOGY 4.51 35.03 12603 48
GENETICS & HEREDITY 4.50 20.02 42825 170
HISTORY 0.61 1.24 5517 100
INFORMATION SCIENCE & LIBRARY SCIENCE 3.38 6.16 8823 85
LAW 1.29 2.36 9059 148
LOGIC 0.51 1.00 1599 19
MATHEMATICS 1.17 2.56 55051 316
MEDICINE, GENERAL & INTERNAL 4.38 20.42 62300 167
MULTIDISCIPLINARY SCIENCES 6.03 13.95 120752 71
NANOSCIENCE & NANOTECHNOLOGY 8.12 12.88 85204 106
ONCOLOGY 5.70 53.73 92083 240
PHYSICS, MULTIDISCIPLINARY 3.83 9.20 42287 78
PSYCHOLOGY, PSYCHOANALYSIS 0.59 1.19 833 12
SOCIAL SCIENCES, INTERDISCIPLINARY 2.25 3.55 12451 108
Table 1: Citation averages $\mu_{s}$ and standard deviations $\sigma_{s}$ for papers in various research fields, which are defined here as categories in the 2020 Journal Citation Reports (JCR) of Clarivate Analytics. The third column shows the number of papers and the fourth column shows the number of journals in each subject. The first entry corresponds to treating all 12173 journals in the 2020 JCR as a single subject.
3 Unlike the Impact Factor, the $\Phi$ index is scale independent and passes the ‘random sample test’
As we showed earlier, the citation average $f(n)$ of a randomly formed journal has a range of values that depend on journal size $n$. Specifically, the range of $f(n)$ values measured from the population mean $\mu$ vary as $\sigma/\sqrt{n}$. We demonstrated this effect with the numerical experiment of Figure 1, where we plotted $f(n)$ vs. $n$ for 20,000 randomly drawn ‘journals’ in the subject of cell biology.
Does the $\Phi$ index pass this ‘random sample test’? If so, then plotting the $\Phi$ index vs. size would yield no size dependence. In Figures 5, 6, 7, and 8, we plot the citation average $f(n)$ (left panels) and $\Phi$ index (right panels) vs. journal size for 20,000 ‘journals’ drawn randomly from papers in the subjects of multidisciplinary sciences, condensed matter physics, psychology, and information science & library science, respectively. Evidently, the range of values of the $\Phi$ index in all these cases shows no dependence with journal size, unlike the citation average. Furthermore, the vast majority of $\Phi$ index values are bounded within the range $[-3,3]$. These results make sense and confirm our expectation from Eq. (6), where the inequality covers 99% of random journals for $k=3$.
Figure 5: Citation average (left) and $\Phi$ index (right) vs. size for a randomly sampled journal in multidisciplinary sciences. The citation average is scale dependent, the $\Phi$ index not.
Figure 6: Citation average (left) and $\Phi$ index (right) vs. size for a randomly sampled journal in condensed matter physics. The citation average is scale dependent, the $\Phi$ index not.
Figure 7: Citation average (left) and $\Phi$ index (right) vs. size for a randomly sampled journal in psychology. The citation average is scale dependent, the $\Phi$ index not.
Figure 8: Citation average (left) and $\Phi$ index (right) vs. size for a randomly sampled journal in information science and library science. The citation average is scale dependent, the $\Phi$ index not.
Figure 9: Citation average (left) and $\Phi$ index (right) vs. size for randomly sampled journals in physics (2020 JCR). The citation average is scale dependent, the $\Phi$ index not. Furthermore, the randomly formed journals have citation averages that range between the theoretically expected upper and lower values from Eq. (3), shown here (for $k=3$) as continuous blue and orange lines, respectively. The $\Phi$ index values are confined within the range [3,3[-3,3[ - 3 , 3, as expected from Eq. (6) for 99% of journals. Shown in red are the citation averages of actual journals in physics.
4 $\Phi$ index for journals
Having established the mathematical foundation and the statistical relevance of the $\Phi$ index, we now proceed to calculate it for journals in several subjects.
4.1 Method & Data
First, for each journal in the 2020 JCR we obtained its individual Journal Citation Report, which contained the 2020 citations to each of its citable papers (i.e., articles and reviews) published in 2018–2019. (143 journals whose individual citation reports would not download due to technical glitches were ignored.) We thus ended up with a master list of 12,173 journals with citation distributions. For each journal in the master list we calculated its citation average, $f$, which approximates the IF and becomes identical to it when there are no “free” or “stray” citations in the numerator (these are citations to front-matter items such as editorials, letters to the editor, commentaries, etc., or citations to the journal without specific reference of volume and page or article number). We will thus use the terms “IF” and “citation average” interchangeably, for simplicity. Collectively, the 12,173 journals in our master list published 3,537,118 papers in 2018–2019, which received 14,550,739 citations in 2020. This is our data set.
4.2 Results I: Scale-independent $\Phi$ index
As a first step, we calculate the $\Phi$ index for all journals in the JCR as if they belong to the same subject.
Do the different rankings make sense? In Tables 2 and 3 we list the top-50 journals ranked by the overall $\Phi$ index and the article-only $\Phi_{A}$ index, respectively; we also list the corresponding citation average rankings ($f$-ranks and $f_{A}$-ranks, respectively), for comparison. We can make the following observations.
(a) The $\Phi$ rankings are sufficiently distinct from the $f$ rankings. For example, 20 of the top-50 journals ranked by $\Phi$ are new entries. Likewise, 123 of the top-500 $\Phi$-ranked journals are new entries.
(b) Importantly, several entries in the top 50 $\Phi$ rankings improved considerably their positions compared to $f$ rankings. Of particular note are the following transitions from rank($f$) to rank($\Phi$):
NATURE COMMUNICATIONS (#196 $\to$ #10),
JOURNAL OF THE AMERICAN CHEMICAL SOCIETY (#173 $\to$ #15),
ANGEWANDTE CHEMIE (#193 $\to$ #16),
ADVANCED FUNCTIONAL MATERIALS (#128 $\to$ #18),
APPLIED CATALYSIS B-ENVIRONMENTAL (#118 $\to$ #19),
NUCLEIC ACIDS RESEARCH (#151 $\to$ #24),
JOURNAL OF MATERIALS CHEMISTRY A (#250 $\to$ #26),
CHEMICAL ENGINEERING JOURNAL (#226 $\to$ #28),
ACS NANO (#170 $\to$ #29),
NANO ENERGY (#131 $\to$ #30),
PNAS (#334 $\to$ #36),
ACS APPLIED MATERIALS & INTERFACES (#453 $\to$ #41), and
RENEWABLE & SUSTAINABLE ENERGY REVIEWS (#189 $\to$ #49).
(c) The journals NATURE, SCIENCE, and CELL improve their rankings and enter the top-10 slots. Intuitively, these results make sense. Indeed, these improved rankings correspond better with the reputation of these three titles as among the most selective journals in the world. Other new entries in the top-10 $\Phi$ ranks are ADVANCED MATERIALS, CHEMICAL SOCIETY REVIEWS, and NATURE COMMUNICATIONS.
Figure 10 shows $\Phi$ rankings plotted against $f$ rankings for all 12173 journals (as if they belong to a unified subject, and with articles and reviews counted together). We note that the two rankings converge at rank #2382, the “inflection” point shown in Fig. 10. At the inflection point, the $\Phi$ index flips sign and the citation average $f$ crosses the global citation average $\mu$. So, to the left of the inflection point we have all 2382 journals for which $f>\mu$ and $\Phi>0$. To the right of the inflection point lie all the 9791 journals for which $f<\mu$ and $\Phi<0$.
How similar are the two rankings? A measure of rank correlation—the similarity of journal orderings when ranked by $f$ vs. $\Phi$—is given by the Kendall $\tau$ coefficient, which takes the value $\tau=0.62$ for the full set of 12173 journals, while $\tau=0.73$ for the 2382 journals of positive $\Phi$. While the two rankings have some similarity, it is clear from the Kendall coefficient that substantial re-ordering occurs, as we also see from Fig. 10.
Figure 10: Left panel: Comparison of $\Phi$ rankings vs. $f$ rankings for all 12173 journals. The convex-shaped “lobe” on the lower left contains the 2382 journals with $\Phi>0$ or $f>\mu$. The large concave-shaped “lobe” on the upper right contains the 9791 journals with $\Phi<0$ or $f<\mu$. A measure of rank correlation between the two rankings is given by the Kendall coefficient $\tau=0.62$. Right panel: Detail of left panel for the 2382 ranks whereby $\Phi>0$ or $f>\mu$. For this set, $\tau=0.73$.
Journal $N_{2Y}$ $f$ $\Phi$ rank($f$) rank($\Phi$)
CA-A CANCER JOURNAL FOR CLINICIANS 47 503.9 274 1 1
NEW ENGLAND JOURNAL OF MEDICINE 649 79.8 154.1 3 2
NATURE 1807 45.4 140.5 21 3
SCIENCE 1556 42.1 119.9 23 4
LANCET 539 68.5 119.6 5 5
ADVANCED MATERIALS 2808 29.6 107.8 51 6
CHEMICAL REVIEWS 431 59.9 92.6 8 7
CHEMICAL SOCIETY REVIEWS 482 53.7 87.1 15 8
CELL 886 39.4 84.1 30 9
NATURE COMMUNICATIONS 10531 14.3 83.8 196 10
NATURE MEDICINE 391 51.5 74.9 16 11
ADVANCED ENERGY MATERIALS 1428 28.3 73 57 12
NATURE REVIEWS MOLECULAR CELL BIOLOGY 90 90.7 65.7 2 13
ENERGY & ENVIRONMENTAL SCIENCE 564 38 64.3 33 14
JOURNAL OF THE AMERICAN CHEMICAL SOCIETY 4924 15.3 62.9 173 15
ANGEWANDTE CHEMIE-INTERNATIONAL EDITION 5567 14.6 62.4 193 16
NATURE ENERGY 204 58 61.5 11 17
ADVANCED FUNCTIONAL MATERIALS 2751 18.1 58.4 128 18
APPLIED CATALYSIS B-ENVIRONMENTAL 2136 19.2 55.8 118 19
JAMA-J AMER MED ASSOC 412 36.5 52.6 39 20
NATURE MATERIALS 308 41.1 51.9 24 21
NATURE BIOTECHNOLOGY 228 45.7 50.2 20 22
NATURE GENETICS 361 37.1 50.1 37 23
NUCLEIC ACIDS RESEARCH 2535 16.4 49.5 151 24
JOULE 339 37.4 49 35 25
JOURNAL OF MATERIALS CHEMISTRY A 5182 12.6 48.9 250 26
LANCET ONCOLOGY 326 38 48.9 32 27
CHEMICAL ENGINEERING JOURNAL 4657 13.1 48.9 226 28
ACS NANO 2670 15.6 47.3 170 29
NANO ENERGY 1835 17.7 46.6 131 30
NATURE NANOTECHNOLOGY 300 37.5 46.2 34 31
CIRCULATION 716 25.5 45.7 67 32
NATURE REVIEWS MATERIALS 91 63.5 45.3 6 33
NATURE REVIEWS MICROBIOLOGY 109 58.4 45.3 10 34
NATURE REVIEWS DRUG DISCOVERY 72 69 44 4 35
PNAS 6620 10.8 43.3 334 36
JOURNAL OF CLINICAL ONCOLOGY 642 25.3 42.8 70 37
NATURE REVIEWS CANCER 95 58.5 42.4 9 38
NATURE CATALYSIS 214 39.9 41.9 28 39
NATURE REVIEWS CLINICAL ONCOLOGY 83 60.5 41.1 7 40
ACS APPLIED MATERIALS & INTERFACES 10072 9.2 40.7 453 41
NATURE PHOTONICS 218 36.9 38.7 38 42
EUROPEAN HEART JOURNAL 519 25.1 38.2 72 43
IMMUNITY 330 30.3 38 47 44
ACS ENERGY LETTERS 723 21.7 37.9 93 45
NATURE REVIEWS IMMUNOLOGY 104 49.8 37.3 19 46
ACCOUNTS OF CHEMICAL RESEARCH 645 22.1 36.6 89 47
LANCET NEUROLOGY 159 40.3 36.5 25 48
RENEWABLE & SUSTAINABLE ENERGY REVIEWS 1801 14.8 36.2 189 49
NATURE REVIEWS GENETICS 95 50.4 36.1 18 50
Table 2: Top 50 journals ranked by their overall $\Phi$ index.
(c) What about different document types? We know that articles and reviews are cited differently (Garfield, 1996; MacRoberts & MacRoberts, 1996), which calls for separate journal rankings for articles and reviews. This concern can be easily accommodated. One can obtain $\Phi$ index rankings for journals that publish only articles or reviews, or a mixture of both. See Table 3, where we list the top 50 ranked journals in terms of their article-only $\Phi_{A}$ index. Compared to Table 2, we note that many journals retain their positions in the top-50 slots, which means that the $\Phi$ and $\Phi_{A}$ indexes are largely correlated—indeed, Kendall $\tau(\text{rank}(\Phi),\text{rank}(\Phi_{A}))=0.91$. But upon closer inspection, we note that review journals are over-represented in the top $\Phi$ ranks of Table 2, which is a clear consequence of the fact that review articles skew the rankings (Miranda & Garcia-Carpintero, 2018). For example, 11 journals or 22% of the top-50 titles in Table 2 publish only reviews. Some notable transitions in the top-50 list from rank($f_{A}$) to rank($\Phi_{A}$) are the following:
JOURNAL OF MATERIALS CHEMISTRY A (#168 $\to$ #27),
SCIENCE ADVANCES (#137 $\to$ #34),
JOURNAL OF CLEANER PRODUCTION (#339 $\to$ #39),
PHYSICAL REVIEW LETTERS (#313 $\to$ #43), and
SCIENCE OF THE TOTAL ENVIRONMENT (#504 $\to$ #48).
As a cautionary tale about data quality and data cleaning, we note in Table 3 the titles NATURE REVIEWS DISEASE PRIMERS and REVIEWS OF MODERN PHYSICS. Both these entries reflect a cataloging error in the Document Type by Clarivate Analytics, since we know that these journals publish only reviews and should thus be absent from article-only $\Phi_{A}$ rankings of journals.
Journal $N_{2Y,A}$ $f_{A}$ $\Phi_{A}$ rank($f_{A}$) rank($\Phi_{A}$)
CA-A CANCER JOURNAL FOR CLINICIANS 35 659.1 321.1 1 1
NEW ENGLAND JOURNAL OF MEDICINE 553 84.7 157.6 3 2
NATURE 1755 44.7 141.9 15 3
SCIENCE 1434 40.2 114.3 18 4
LANCET 383 65.8 100.5 4 5
ADVANCED MATERIALS 2450 27.2 96.1 42 6
NATURE COMMUNICATIONS 10471 14.2 88.2 130 7
CELL 815 37.5 79.6 20 8
NATURE MEDICINE 373 48.6 71.8 11 9
ADVANCED ENERGY MATERIALS 1296 25.9 65.8 46 10
JOURNAL OF THE AMERICAN CHEMICAL SOCIETY 4848 15 64.9 116 11
ANGEWANDTE CHEMIE-INTERNATIONAL EDITION 5260 13.4 57.9 143 12
APPLIED CATALYSIS B-ENVIRONMENTAL 2094 18.8 56.9 79 13
ADVANCED FUNCTIONAL MATERIALS 2591 16.9 55.4 89 14
ENERGY & ENVIRONMENTAL SCIENCE 494 32.9 53.6 29 15
NUCLEIC ACIDS RESEARCH 2534 16.4 52.6 92 16
NATURE GENETICS 361 37.1 52.4 22 17
LANCET ONCOLOGY 255 43.3 52.2 16 18
NATURE BIOTECHNOLOGY 218 45.4 50.9 13 19
NATURE ENERGY 174 50.1 50.6 10 20
ACS NANO 2632 15.4 49.3 113 21
CHEMICAL ENGINEERING JOURNAL 4518 12.4 48.2 159 22
NATURE MATERIALS 294 37.1 47.3 21 23
PNAS 6619 10.8 47.1 204 24
CIRCULATION 690 25.3 46.7 49 25
NANO ENERGY 1771 17.2 46.7 88 26
JOURNAL OF MATERIALS CHEMISTRY A 4896 11.8 46.5 168 27
JOURNAL OF CLINICAL ONCOLOGY 581 26.8 45.9 44 28
NATURE NANOTECHNOLOGY 286 36.4 45.7 23 29
ACS APPLIED MATERIALS & INTERFACES 10032 9.1 44.1 299 30
JOULE 269 35.5 43.1 24 31
JAMA-J AMER MED ASSOC 304 31.9 40.6 33 32
ACS ENERGY LETTERS 701 21 37.8 68 33
SCIENCE ADVANCES 2103 13.5 37 137 34
NATURE CATALYSIS 193 35.5 36.5 25 35
NATURE REVIEWS DISEASE PRIMERS 82 51.2 35.6 9 36
EUROPEAN HEART JOURNAL 449 23.8 35.2 54 37
LANCET NEUROLOGY 101 45.2 34.4 14 38
JOURNAL OF CLEANER PRODUCTION 7409 8.6 34.4 339 39
ACS CATALYSIS 2317 12.3 34.1 161 40
NATURE PHOTONICS 202 31.9 33.1 32 41
JAMA ONCOLOGY 280 27.2 32.4 43 42
PHYSICAL REVIEW LETTERS 5425 8.9 31.4 313 43
REVIEWS OF MODERN PHYSICS 43 60.1 30.6 6 44
IMMUNITY 281 25.6 30.3 47 45
CANCER CELL 216 28.2 29.8 40 46
J AMER COLL CARDIOL 610 18.2 29.5 82 47
SCIENCE OF THE TOTAL ENVIRONMENT 10054 7.3 29.3 504 48
NATURE METHODS 289 24.3 28.9 52 49
NATURE NEUROSCIENCE 339 22.7 28.9 61 50
Table 3: Top 50 journals ranked by their article-only $\Phi_{A}$ index.
4.3 Results II: Scale-independent and subject-specific $\Phi$ index for monodisciplinary journals
Subject category $\mu_{s}^{art}$ $\mu_{s}^{rev}$ $\sigma_{s}^{art}$ $\sigma_{s}^{rev}$ $N_{s}^{art}$ $N_{s}^{rev}$ Jnls
AGRICULTURE, DAIRY & ANIMAL SCIENCE 2.08 5.84 2.64 8.69 17089 759 62
PSYCHIATRY 3.45 7.17 5.37 9.89 38282 5186 217
CLINICAL NEUROLOGY 3.5 6.22 7.91 13.57 51637 9751 208
MEDICINE, GENERAL & INTERNAL 4.01 5.97 20.68 19.21 50457 11843 167
PARASITOLOGY 3.31 6.44 5.15 11.23 11195 1028 37
AGRICULTURAL ENGINEERING 5.59 17.32 6.16 17.21 7917 295 14
GEOCHEMISTRY & GEOPHYSICS 3.13 8.59 4.32 13.57 23175 389 79
MATERIALS SCIENCE, CHAR. & TESTING 2.79 6.2 4.48 10.45 6621 117 30
ALLERGY 4.72 6.42 7.57 8.85 4225 1234 28
CRITICAL CARE MEDICINE 4.63 6.4 10.74 9.22 8584 1568 36
SOCIAL SCIENCES, BIOMEDICAL 2.71 6.08 3.89 6.84 6509 313 44
ENVIRONMENTAL SCIENCES 4.64 12.27 6.77 19.15 145587 6773 273
ENGINEERING, PETROLEUM 2.58 5.41 3.51 6.91 4804 170 17
PHYSICS, ATOMIC, MOLECULAR & CHEMICAL 3.5 10.13 5.05 14.29 33278 325 34
ENERGY & FUELS 6.56 18.45 9.89 25.91 81064 4835 108
GENETICS & HEREDITY 4.07 8.5 20.54 13.89 38614 4211 170
LIMNOLOGY 2.99 5.81 3.79 9.56 4084 79 21
CELL BIOLOGY 6.41 11.95 12.49 24.06 52345 8786 191
PSYCHOLOGY, DEVELOPMENTAL 3.17 7.16 3.97 9.51 11390 609 74
COMPUTER SCIENCE, INTERDISC APPL 3.93 9.56 8.3 16.65 31545 799 111
LINGUISTICS 1.5 3.4 2.69 5.48 11878 283 192
FISHERIES 2.6 6.78 3.47 9.17 12564 473 54
DEVELOPMENT STUDIES 3.04 5.87 4.9 7.67 4529 135 42
HOSPITALITY, LEISURE, SPORT & TOURISM 4.84 11.59 7.56 17.46 6773 307 58
SOCIAL ISSUES 2.11 4.36 3.02 7.52 3736 133 44
PSYCHOLOGY, MATHEMATICAL 3.73 5.77 13.76 8.62 1338 122 13
FAMILY STUDIES 2.43 5.34 3.13 8.79 6199 290 46
CHEMISTRY, PHYSICAL 6.75 20.22 10.74 28.11 144909 4799 150
MATERIALS SCIENCE, BIOMATERIALS 5.31 13.53 6.14 15.03 16647 1140 41
MATERIALS SCIENCE, MULTIDISCIPLINARY 5.37 18.13 9.17 27.66 240166 8437 327
Table 4: Global citation averages, standard deviations, and document numbers for articles and reviews in 30 subject categories. The number of journals in each category is listed in the last column.
So far, we have ranked $\Phi$ indexes of all journals as if they belong to a unified subject, with a global citation average $\mu$ and and a global citation standard deviation $\sigma$. Of course, different subjects have different citation practices, which one needs to account for in order to meaningfully compare citation averages across subjects. As we mentioned in §2, we can readily apply the $\Phi$ index methodology to this task. The different citation practices of each subject category $s$ are simply “coded” in the numbers ($\mu_{s}$, $\sigma_{s}$), which are used to calculate a journal’s $\Phi_{s}$ index, as shown in Eq. (11).
Therefore, for monodisciplinary journals that publish in a single academic discipline, this is a straightforward step to take. We can calculate and compare $\Phi_{s}$ indexes across subjects, because the $\Phi$ index is uniquely suited to this task, in the same way that we use $z$ scores to compare athletes’ performances across different sports.
Likewise, we can readily calculate the composite $\Phi_{s}$ index of a journal whose entirety is allocated to two or more subjects. In this case, all papers published by the journal belong to each subjects. For example, take NATURE PHOTONICS. All of its papers are concurrently allocated by Clarivate in two subject categories, OPTICS and PHYSICS, APPLIED. So, the composite $\Phi_{s}$ index of the journal is simply the average of its $\Phi_{s}$ indexes in each subject. And since NATURE PHOTONICS has $\Phi_{s}=68.3$ in OPTICS and $\Phi_{s}=42.5$ in PHYSICS, APPLIED, its composite $\Phi_{s}=55.4$.
Journal $n_{A}$ $f_{A}$ $\Phi_{A}$ $\Phi_{s,A}$ rank($f_{A}$) rank($\Phi_{A}$) rank($\Phi_{s,A}$)
NATURE 1755 44.7 141.9 121 15 3 1
ADVANCED MATERIALS 2450 27.2 96.1 103.3 42 6 2
SCIENCE 1434 40.2 114.3 96.8 18 4 3
NEW ENGLAND JOURNAL OF MEDICINE 553 84.7 157.6 91.7 3 2 4
NATURE MEDICINE 373 48.6 71.8 73.2 11 9 5
ADVANCED ENERGY MATERIALS 1296 25.9 65.8 72.5 46 10 6
ENERGY & ENVIRONMENTAL SCIENCE 494 32.9 53.6 72 29 15 7
CA-A CANCER JOURNAL FOR CLINICIANS 35 659.1 321.1 67.1 1 1 8
NATURE BIOTECHNOLOGY 218 45.4 50.9 65.4 13 19 9
NATURE NEUROSCIENCE 339 22.7 28.9 63.2 61 50 10
NATURE COMMUNICATIONS 10471 14.2 88.2 63.1 128 7 11
JOURNAL OF THE AMER CHEM SOC 4848 15 64.9 63.1 114 11 12
NATURE ENERGY 174 50.1 50.6 61.2 10 20 13
CELL 815 37.5 79.6 60.8 20 8 14
APPLIED CATALYSIS B-ENVIRONMENTAL 2094 18.8 56.9 60.6 78 13 15
LANCET 383 65.8 100.5 58.5 4 5 16
NATURE PHOTONICS 202 31.9 33.1 55.4 32 41 17
ADVANCED FUNCTIONAL MATERIALS 2591 16.9 55.4 54.8 87 14 18
ANGEWANDTE CHEMIE-INTERN EDIT 5260 13.4 57.9 54.1 141 12 19
NATURE MATERIALS 294 37.1 47.3 54.1 21 23 20
LANCET NEUROLOGY 101 45.2 34.4 52.9 14 38 21
NATURE NANOTECHNOLOGY 286 36.4 45.7 50.3 23 29 22
IMMUNITY 281 25.6 30.3 49.5 47 45 23
IEEE COMMUN SURV AND TUTOR 236 24.6 26.5 48.7 51 58 24
JOULE 269 35.5 43.1 48.6 24 31 25
NANO ENERGY 1771 17.2 46.7 47.2 86 26 26
NEURON 582 14.6 21.7 47.1 123 85 27
CHEMICAL ENGINEERING JOURNAL 4518 12.4 48.2 46.6 157 22 28
PHYSICAL REVIEW LETTERS 5425 8.9 31.4 46.2 307 43 29
ACS NANO 2632 15.4 49.3 45.3 111 21 30
MMWR SURVEILLANCE SUMMARIES 26 57.8 22.8 44.5 7 75 31
REVIEWS OF MODERN PHYSICS 43 60.1 30.6 44 6 44 32
NATURE CLIMATE CHANGE 268 20.9 23.2 43.4 70 70 33
CARBOHYDRATE POLYMERS 2396 8.7 19.8 42 326 98 34
ACS ENERGY LETTERS 701 21 37.8 41.7 68 33 35
MMWR-MORBID MORTAL WEEKLY REP 456 14.9 19.7 41.1 118 102 36
NATURE PLANTS 215 14.6 13.1 40.7 124 163 37
LANCET INFECTIOUS DISEASES 198 22.8 22.2 40.6 59 80 38
JOURNAL OF MATERIALS CHEMISTRY A 4896 11.8 46.5 39.7 166 27 39
SCIENCE OF THE TOTAL ENVIRONMENT 10054 7.3 29.3 39.7 495 48 40
Table 5: Top 40 journals ranked by subject $\Phi$ index (articles only), $\Phi_{s,A}$. $n_{A}$ is each journal’s size. Shown for comparison are the rankings by citation average $f{A}$ and $\Phi_{A}$.
Once we have subject-specific $\Phi$ indexes (denoted $\Phi_{s}$), we can compare journals in different subjects. In Table 6, we list the top 40 journals ranked by their article-only $\Phi_{s}$ index, i.e., $\Phi_{s,A}$. This ranking corrects for subject disparities, in addition to scale disparities that are already corrected by use of the general $\Phi$ index. Because journals in less-cited subjects are compared to their in-subject journals (via use of $\mu_{s}$ and $\sigma_{s}$), their citation averages are subject-standardized, which allows them the opportunity to rise up in rankings, if they are sufficiently cited within their subject.
With the decrease of subject disparity in standardized citation averages we expect journals from a wider range of subjects in the higher $\Phi_{s}$ ranks. And indeed this is what happens. For example, the top 1000 slots of $\Phi_{s,A}$ rankings include journals from nearly all subjects (220 subject categories out of 229 JCR categories, or 96%). Compare this with 180 categories in the top-1000 $\Phi_{A}$ rankings (79%), or 178 categories in the top-1000 $f_{A}$ rankings (78%).
In particular, the top 1000 $\Phi_{s,A}$ ranks include more journals from mathematics, sociology, entomology, family studies, agronomy, history, linguistics, and law, which are some of the underrepresented categories when we rank by citation average, $f_{A}$. At the same time, the top 1000 $\Phi_{s,A}$ ranks list fewer journals from materials science, management, oncology, business, cell biology, biochemistry, chemistry. An indicative list of some new entries in the top 1000 slots (compared to when we ranked by citation average, $f_{A}$) is shown in Table 6.
Subject Title
MATHEMATICS ANNALS OF MATHEMATICS
INVENTIONES MATHEMATICAE
JOURNAL OF THE AMERICAN MATHEMATICAL SOCIETY
SOCIOLOGY AMERICAN JOURNAL OF SOCIOLOGY
JOURNAL OF MARRIAGE AND FAMILY
WORK EMPLOYMENT AND SOCIETY
ENTOMOLOGY INSECT SYSTEMATICS AND DIVERSITY
INSECT BIOCHEMISTRY AND MOLECULAR BIOLOGY
PEST MANAGEMENT SCIENCE
FAMILY STUDIES JOURNAL OF MARRIAGE AND FAMILY
CHILD ABUSE & NEGLECT
AGRONOMY FIELD CROPS RESEARCH
EUROPEAN JOURNAL OF AGRONOMY
INDUSTRIAL CROPS AND PRODUCTS
HISTORY PAST & PRESENT
HISTORY & PHILOSOPHY OF SCIENCE BRITISH JOURNAL FOR THE PHILOSOPHY OF SCIENCE
SOCIAL STUDIES OF SCIENCE
LINGUISTICS APPLIED LINGUISTICS
STUDIES IN SECOND LANGUAGE ACQUISITION
MODERN LANGUAGE JOURNAL
LAW HARVARD LAW REVIEW
STANFORD LAW REVIEW
YALE LAW JOURNAL
Table 6: An indicative list of some new entries in the top 1000 ranks when we rank by $\Phi_{s,A}$ (subject-specific $\Phi$ index) as opposed to ranking by $f_{A}$ (citation average).
4.3.1 Cautionary note: Over-analyzing subject categories
A cautionary note is warranted here. The categorization of journals into increasingly more specific categories may give a misleading sense of accuracy while in fact it can lead to flawed conclusions when we calculate (and standardize) subject-specific citation averages. This is because when we standardize within a subject, the choice of journals that constitute a subject matters a lot, because it affects the subject citation average $\mu_{s}$, and therefore the sign of the $\Phi_{s}$ index. With increasing categorization the number of journals in a subject decreases and $\Phi_{s}$ indexes become more volatile, because the inclusion (or omission) of even a single journal in the subject list can throw another journal’s $\Phi_{s}$ index up or down the ranks as its citation average $f_{A}$ crosses the subject average $\mu_{s}$. As we mentioned earlier in §2.1, the choice of which journals constitute a subject is critical.
For example, to calculate subject-specific $\Phi$ indexes of physics journals, we can use the Clarivate lists of physics journals. Clarivate lists physics journals in 8 subcategories, which we can group together as one category “Physics.” In Table 7, we list the subject specific citation means and standard deviations for each subcategory, as well as for all subcategories together.
Field (subject) $\mu_{s}$ $\sigma_{s}$ no. papers no. journals
PHYSICS 4.1 9.5 302089 413
PHYSICS, MATHEMATICAL 2.3 4.8 20953 53
PHYSICS, NUCLEAR 2.5 4.7 11049 17
PHYSICS, FLUIDS & PLASMAS 2.6 3.5 18892 30
PHYSICS, ATOMIC, MOLECULAR & CHEMICAL 3.6 5.3 33603 34
PHYSICS, MULTIDISCIPLINARY 3.8 9.2 42287 78
PHYSICS, PARTICLES & FIELDS 4.1 14.6 25240 26
PHYSICS, APPLIED 5 10.4 148133 152
PHYSICS, CONDENSED MATTER 6.3 12.7 69916 66
Table 7: Citation average, $\mu_{s}$, and standard deviation, $\sigma_{s}$, for physics papers (articles & reviews) in various subject categories. Note that some journals are listed in more than one subject categories.
4.4 Scale-independent and subject-specific $\Phi$ index for multidisciplinary journals
How do we calculate subject-specific $\Phi$ indexes for cross-disciplinary journals that publish in various subjects, such as Nature, Science, PLOS ONE, Scientific Reports, etc? Here, while every paper published in the journal may belong to a single subject—even though a paper often belongs to more than one subjects, e.g., a physics paper may be also classified as an optics paper—the journal itself publishes papers from various subjects. So, calculation of a journal-specific ($\mu_{s},\sigma_{s}$) is not straightforward, and we need to exercise care.
There are two approaches to this problem: (a) Treat a multidisciplinary journal as a composite of single-subject ‘sub-journals’ and calculate a $\Phi_{s}$ index for each sub-journal. This way, we would end up with as many $\Phi_{s}$ indexes for a journal as the number of subjects it covers. (b) Calculate a composite $\Phi$ index for all the subjects that comprise the journal, and use that single metric to describe the multidisciplinary journal.
Both approaches have pros and cons. In the former approach it may seem confusing that a journal such as Science may have several $\Phi_{s}$ indexes. However, this situation is not unlike recognizing that, say, a university with different departments has a different ranking score in each area of activity, so it can be higher ranked in some areas than others. To calculate $\Phi_{s}$ indexes for multidisciplinary journals along these lines, we need to allocate each paper to at least one research subject. This introduces the complication of having to determine subjects on the article level.
In the latter approach, the task is to integrate various subject-specific $\Phi_{s}$ indexes to produce a single, composite, $\Phi$ index for the whole journal. While this approach may seem simpler at a technical level—all we need is build a composite that makes statistical sense, using, e.g., weighted averages—its conceptual challenge is the use of a single metric to describe the citation performance of disparate fields in the same journal. On the other hand, we can envision practical situations where users may find a composite $\Phi$ index advantageous for a quick and rough impression, in the same vein that an applicant may be interested in a university’s overall position in academic rankings (not in specific subjects), a journalist in a country’s overall health index score (not in specific categories such as air pollution or infant mortality), and an economist in a country’s overall unemployment rate (not in age- or sector-specific rates).
In future work, we will explore the $\Phi$ index of multidisciplinary journals further.
5 $\Phi$ index for other citation averages
We have seen how the $\Phi$ index arises naturally from the Central Limit Theorem, as the distance of the citation average from the mean, divided (rescaled) by the sample standard deviation. Although we have applied the $\Phi$ index to journals so far, there is nothing specific to journals in how the $\Phi$ index was developed. We can thus readily adapt the $\Phi$ index methodology to standardize for scale and subject and compare the citation averages of individual researchers, research teams, universities, and countries.
6 Conclusions and outlook
In this work, we introduced the $\Phi$ index, a scale- and subject-independent citation average indicator. We explained how the $\Phi$ index is based on the Central Limit Theorem of Statistics, via standardization of citation averages in a process that is similar to calculating the $z$ score and that has an elegant geometric interpretation. Thus, the $\Phi$ index of a sample of $n$ scientific papers is the distance of the sample’s citation average $f$ from the subject citation mean $\mu$, measured (normalized) in units of the sample’s standard deviation $\sigma_{n}=\sigma/\sqrt{n}$.
We described a simple numerical check for testing any citation indicator for scale- and subject-independence. We then applied the $\Phi$ index to scientific journals. We compared $\Phi$-index with citation-average ($f$) rankings for journals using data from the 2020 Journal Citation Reports. We showed how the scale standardization of the $\Phi$ index corrects for the scale bias in $f$ rankings, and how the subject standardization allows journals from a more diverse range of subjects (such as mathematics, sociology, entomology, family studies, agronomy, history, linguistics, and law) in the top 1000 slots of $\Phi$ rankings. The $\Phi$-index methodology is mathematically sound, statistically robust, easy to implement, and can be extended to other populations such as researchers, universities, or countries.
Future work could involve the application of the $\Phi$ index to individual researchers or universities, or the comparison of the $\Phi$ index with other subject-normalized indicators for journals. It would also be interesting to study whether and to what extent the $\Phi$ index is vulnerable to gaming. Another project would be to study the extent to which the $\Phi$ index is indicative of the “high-end” of the underlying citation distributions, i.e., of papers that are highly cited, or even highlighted by the reputable science press. A preliminary study (Antonoyiannakis, 2023) showed that the $\Phi$ index is also a better descriptor of a journal’s potential for capturing highly cited papers (at the 99th percentile in their academic field per publication year) than the IF.
Acknowledgements
I am grateful to Jerry I. Dadap for stimulating discussions, and to Richard Osgood Jr. and Irving Herman for encouragement and hospitality. This work uses data, accessed through Columbia University, from the Web of Science and Journal Citation Reports (2020).
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Author Contributions
Manolis Antonoyiannakis: Conceptualization, data curation, formal analysis, methodology, writing.